\vspace{0.5cm}


\begin{thebibliography}{99}

\bibitem[B1]{B1} S. Bloch, Motives associated to graphs, preprint, 20 Oct. 2006

\bibitem[CL]{CL} F. Chung, R. Langlands, A combinatorial Laplacian with vertex weights, J. of Combinatorial theory, (A), 75,(1996), 316-327


\bibitem[ITW]{ITW} E. Izadi, CS. Tama, J. Wang, The primitive cohomology of the Theta divisor of an abelian 5-fold, Arxiv e-print 1311.6212, 


\bibitem[JS2]{JS2} J. Steenbrink: Mixed Hodge structure on the vanishing cohomology. In: P. Holm (ed.): Real and complex Singularities. Oslo 1976. Sijthoff-Noordhoff, 
1977, pp. 525-563

\bibitem[JS3]{JS3} J. Steenbrink: Mixed Hodge structures associated with isolated singularities. Proc. Symp. Pure Math. 40, Part 2 (1983) 513-536


\bibitem[S]{S} J. P. Serre, Algebraic groups and class fields, Graduate Texts in Mathematics, Vol. 117 , translation of the French edition; Groupes algebriques et corps de classes. Paris: Hermann, 1975

\bibitem[SP]{SP}  Steenbrink J. , Peters C. , Mixed Hodge structures, A series of modern surveys in mathematics. Springer verlag publications, Vol 52, 2007



\end{thebibliography}

\documentclass[12pt,twoside]{amsart}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amscd}
\usepackage{xypic}
\xyoption{all}
\title[Motive of Theta divisor I]{Motive of Theta divisor I}

\author{Mohammad Reza Rahmati}
\thanks{}
\address{Centro de Investigacion en Matematicas, CIMAT
\hfill\break 
\hfill\break \\
\hfill\break }
\email{mrahmati@cimat.mx}

\newcommand{\comments}[1]{}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{proposition}[theorem]{Proposition}

\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{definition}[theorem]{Definition}
\newtheorem{remark}[theorem]{Remark}

\keywords{Mixed Hodge structure, Generalized Jacobian, Riemann-Roch Theorem, Motives, Periods, Algebraic group}

\subjclass{14-xx}

\begin{document}

\begin{abstract}
It is a well known fact that the Theta divisor on the Jacobian of a non-singular curve is a determinantal variety, i.e. is defined by the zero set of a determinant.
A less known is for the generalized Jacobian of a singular curve. We describe that the method of proof of the first statement can be generalized to the case of singular curves. We continue to study the mixed Hodge structure 
of the generalized Theta divisor using standard techniques in Hodge theory.

\end{abstract}

\maketitle


\section*{Introduction}

Long time, one probably looks for a valuable question in the theory of mixed Hodge structures which involves a rich algebraic geometry. Despite of high lighted questions such as Hodge conjecture or Standard conjectures which look so hard to be solved, perhaps the question of studying the mixed Hodge structure 
of Theta divisor on the Jacobian of curves is one of the very complicated questions. I got acquainted with this question in a lecture by S. Bloch in Madrid at ICMAT. His lectures on extensions of Hodge structures began almost with this concept. The idea that the method can be generalized to the singular curves reached to me at the same time. After searching a little on the question I found, because of very singularities of Theta, the MHS or what we call here the motive of Theta seems to be so complicated object to describe. 

\vspace{0.5cm}

In this note I have tried to interface the idea with my knowledge on algebraic curves which I learnt from the book 
'Algebraic groups and class fields', by J. P. Serre. We target to describe the MHS on the cohomologies of generalized $\Theta$-divisors on (compactification of) the generalized Jacobians of Singular curves. Our approach uses modulus on normal projective curves and the theory of Riemann-Roch for singular curves. The aforementioned reference provides an almost complete picture of the theory of generalized Jacobians both over $\mathbb{C}$ and positive characteristic. We work only in char $0$ case, i.e. over the complex numbers. Theory of modulus on curves provides a simple way to describe the generalized Jacobians and their relation with the Jacobian in the smooth case as extensions of algebraic groups. 

\vspace{0.5cm} 

\section{singular Curves}

\vspace{0.5cm}

The Riemann-Roch theorem in complex dimension 1, can be easily generalized to singular projective curves, using theory of modulus. The proof is a modification of that in smooth case, by careful study of linear series on curves.
In order to fix the idea, let $X$ be a complete, irreducible, non-singular curve and let $\mathfrak{m}=\sum n_p.P$ be a modulus on $X$. let $S$ be the support of $\mathfrak{m}$. Set $X^{\prime}=(X-S) \cup Q$, where $Q$ is one point, and put $\mathcal{O}^{\prime}_Q:=\cap_{P \to Q}\mathcal{O}_P$. Define the sheaf $\mathcal{O}^{\prime}$ by setting it to be equal to $\mathcal{O}$ for the points not in $S$. It is an easy exercise in algebraic geometry to show that the sheaf $\mathcal{O}^{\prime}$ endows $X^{\prime}$ with a structure of algebraic curve such that $X$ is its normalization and its singular points are $Q$. One shows that any singular curve is obtained in this way.

\vspace{0.5cm}

\noindent
The basic example is $\mathfrak{m}=2P$ where one gets an ordinary cusp, and another when $\mathfrak{m}=P_1+P_2$ where one obtains ordinary double point, etc....

\vspace{0.5cm}

The Riemann-Roch theorem for singular curves is written as

\[ l'(D)-i'(D)=\deg(D)+1-\pi \]  

\[ l'(D):=\dim H^0(X',L'(D)), \ i'(D):=\dim H^1(X',L'(D)) \]

\vspace{0.5cm}

\noindent
where $\pi$ is the arithmetic genus. The sheaf $L'(D)$ is the same as $L(D)$ on $X-S$ and equals $\mathcal{O}_Q'$ on $S$. We have the short exact sequence

\[ 0 \to \mathcal{O}' \to \mathcal{O} \to \mathcal{O}/\mathcal{O}' \to 0 
\]

\vspace{0.5cm}

\noindent
The sheaf $\mathcal{O}/\mathcal{O}'$ is concentrated on $S$ and its annihilator ideal is called the conductor of $\mathcal{O}$ into $\mathcal{O}'$. Thus,

\[ 
\chi(X',\mathcal{O}/\mathcal{O}')=\dim H^0(X',\mathcal{O}/\mathcal{O}')=\sum_{Q \in S} \dim \mathcal{O}_Q/\mathcal{O}'_Q=: \delta
\]

\vspace{0.5cm}

\noindent
One can prove that $\pi=g +\delta$ is equal to the arithmetic genus, and in case the curve is defined by an equation of degree $d$ in a projective space $\pi=1+d(d-3)/2, \ g=1-\delta +d(d-3)/2$, the classical Plucker formulas,\cite{S}. 

\vspace{0.5cm}
 
\section{Generalized Jacobians}

\vspace{0.5cm}

Assume $X$ is a projective irreducible algebraic curve. If $f:X \to \mathcal{J}$ be a rational map the the set of points where $f$ is not regular is a finite set $S$. When $\mathcal{J}$ is abelian, then $S=\emptyset$, and $f(D)=0$ if and only if $D$ is the divisor of a rational function. For $f$ regular away from $S$, there can be find a modulus 
$\mathfrak{m} $ with support $S$, such that $f(D)=0$ iff $D=(h)$ with $h \equiv 1 \mod \mathfrak{m}$ (i.e. $v_P(1-h) >n_P$, where $\mathfrak{m}=\sum n_PP$), \cite{S}, Chap. V.

\vspace{0.5cm}

\begin{theorem} (\cite{S} Chap. V, page 88)
For every modulus $\mathfrak{m}$, there exists a commutative algebraic group $J_{\mathfrak{m}}$ and a rational map $\phi_{\mathfrak{m}}:X \to J_{\mathfrak{m}}$ which is universal for all the rational $f:X \to G$ regular away from $S$. This means $f$ factors through $\phi_{\mathfrak{m}}$ as $f=\theta \circ \phi_{\mathfrak{m}}$ for a unique rational homomorphism $\theta:J_{\mathfrak{m}} \to G$. 
\end{theorem}

\vspace{0.5cm}

\noindent
The map $\phi_{\mathfrak{m}}$ defines by extensions to divisor classes a bijection from  $C_{\mathfrak{m}}^0$: the group of divisors of degree zero which are prime to $S$ modulo those of the form $(f)$ where $f \equiv 1 \mod \mathfrak{m}$ to $J_{\mathfrak{m}}$. One has the following exact sequence 

\[ 0 \to R_{\mathfrak{m}}/\mathbb{G}_m \to C_{\mathfrak{m}}^0 \to J \to 0 \] 

\vspace{0.5cm}

\noindent
where $R_{\mathfrak{m}} = \prod_{P \in S} U_P/U_P^{(n_P)}$, and $U_P^{(n)}:=\{f ; v_P(1-f)>n\}$. Each of the groups 
$U/U^{(n)}$ in the product are isomorphic to a product of $\mathbb{G}_m$ with a unipotent group of upper triangular matrices, \cite{S} page 94. 

\vspace{0.5cm}

\noindent
The algebraic group $J_{\mathfrak{m}}$ is called the generalized Jacobian of the curve $X$, and the case $\mathfrak{m}=0$ corresponds to the usual Jacobian $J$. By standard arguments one shows that for $\mathfrak{m} \geq \mathfrak{m}'$ there is a unique surjective separable homomorphism $F:J_{\mathfrak{m}} \to J_{\mathfrak{m}}'$, with connected kernel such that $\phi_{\mathfrak{m}}'=F \circ \phi_{\mathfrak{m}}$. 

\vspace{0.5cm}

\begin{theorem} (\cite{S} page 101)
\[ J_{\mathfrak{m}} \cong \Omega(-\mathfrak{m})^{\vee}/H_1(X-S,\mathbb{Z}) \]
\end{theorem}

\vspace{0.5cm}

The Abel theorem on the Jacobian of non-singular curves asserts that $Pic^0(X) \cong J(X)$, where $Pic^0(X)$ stands for group of zero divisors modulo rational equivalence. This isomorphism is given by integration along paths whose boundray is the given $0$-cycle. Because $J(X)$ is an abelian group, the universal map $\phi:X \to J(X)$, provides a map $sym^n(X) \to J(X)$ whose fibers are projective spaces. For $n=g$ the genus of the curve $X$ this map is surjective and for $n=g-1$ its image is the $\Theta$-divisor on $J(X)$. 

\vspace{0.3cm}

\noindent
We consider $X \times J$ with two projections $pr_1, pr_2$. The universal Poincare line bundle $\wp$ on $X \times J$ is a line bundle s.t. $\wp \mid_{X \times j} \cong \mathcal{O}_X(D(j))$ where $D(j)$ is a divisor class of degree $0$ via Abel theorem. Fix a line bundle $L_0$ a line bundle of degree $g-1$ on $X$. Set $\mathcal{L}:=pr_1^*L_0 \otimes \wp$. Let $D=x_1+...+x_g$ be a divisor of degree $g$ on $X$. Then $D \times J$ is a divisor on $X \times J$ and $\mathcal{O}(D \times J)=pr_1^*\mathcal{O}_X(D)$. Moreover we have $\mathcal{L}(D)=\mathcal{L} \otimes pr_1^*(\mathcal{O}_X(D))$. The following theorem is well-known.

\vspace{0.5cm}

\begin{theorem}
The $\Theta$-divisor is a determinantal subvariety of $J$, i.e is given by the zero set of a determinant function. 
\end{theorem}

\begin{proof}

This can be obtained by applying the Leray cohomological fiber functor $R^{\bullet}pr_{2,*}$ to the short exact sequence of sheaves on $X \times J$,

\begin{equation}
0 \to \mathcal{L} \to \mathcal{L}(D) \to \mathcal{L}(D)/\mathcal{L} \to 0
\end{equation}

\vspace{0.5cm}

\noindent
In the derived functor long exact sequence one obtains 

\begin{equation}
0 \to pr_{2*}\mathcal{L} \to  pr_{2*}\mathcal{L}(D) \to pr_{2*}(\mathcal{L}(D)/\mathcal{L}) \to R^1pr_{2*}\mathcal{L}  \to R^1pr_{2*}\mathcal{L}(D)
\end{equation}

\vspace{0.5cm}

\noindent
We have $pr_{2*}\mathcal{L}=R^1pr_{2*}\mathcal{L}(D)=0$ by Riemann-Roch. Thus 

\begin{equation}
0 \to  pr_{2*}\mathcal{L}(D) \to pr_{2*}(\mathcal{L}(D)/\mathcal{L}) \to R^1pr_{2*}\mathcal{L}  \to 0
\end{equation}

\vspace{0.5cm}

\noindent
where the last sheaf has support the $\Theta$-divisor. Therefore the determinant of the first homomorphism defines $\Theta$. 
\end{proof}

\vspace{0.5cm}

\vspace{0.5cm}
In case of singular curves the canonical map $\phi_{\mathfrak{m}}:X \to J_{\mathfrak{m}}$ by extension over divisors defines a homomorphism from the group of divisors prime to $S$ onto the group $J_{\mathfrak{m}}$. Its kernel is formed by divisors $\mathfrak{m}$-equivalent to an integral multiple of a fixed point $P_0$ in $X'$. 

\vspace{0.5cm}

\noindent
Similar to the non-singular case the map $sym^{\pi} (X) \to J_{\mathfrak{m}}$ is surjective and the image of $sym^{\pi-1} (X) \to J_{\mathfrak{m}}$ is the locus of generalized $\Theta$-divisor. Thus we may apply the argument in the theorem to $X \times J_{\mathfrak{m}}$ and work with divisors prime to $S$ to obtain 

\vspace{0.5cm}

\begin{theorem}
The generalized $\Theta_{\mathfrak{m}}$ is deteminantal on $J_{\mathfrak{m}}$.
\end{theorem}

\vspace{0.5cm}

\section{Motive of Theta divisor}

\vspace{0.5cm}

The motive associated to the $\Theta_{\mathfrak{m}}$ is mainly described by the mixed Hodge structure of $H^{\pi}(J_{\mathfrak{m}} \setminus \Theta_{\mathfrak{m}})$. From now on set $J:=J_{\mathfrak{m}}, \ \Theta:=\Theta_{\mathfrak{m}}$. According to what was said, The $\Theta$-divisor is the singular fiber in a degenerate family defined by a determinant function. The mixed Hodge structure in such a family is determined by J. Steenbrink which uses the method of resolution of singularities. 

\vspace{0.3cm}

\begin{equation}
\begin{CD}
J_{\infty} @>>> U @>>> J @<<< D \\
@Vf_{\infty}VV @VVfV @VVfV @VVV\\
H @>e>> \Delta^* @>>> \Delta @<<< 0 
\end{CD}
\end{equation}

\vspace{0.5cm}

\noindent 
as in 2.9, where $D=\bigcup_i^m D_i$ is a normal crossing divisor. 

\vspace{0.5cm}

\begin{lemma} \cite{JS3}
The spectral sequence of the double complex 

\[ A_D^{pq}:=(a_{q+1})_*\Omega_{D^{(q+1)}}^p, \   \ D^{(q+1)}= \amalg D_{i_1} \cap ... \cap D_{i_q}  \] 

\vspace{0.3cm}

\noindent
with horizontal arrows to be the de Rham differentials, and vertical arrows $\sum_j (-1)^{q+j}\delta_j$, induced by the inclusions $ \amalg D_{i_1} \cap ... \cap D_{i_q} \hookrightarrow \amalg D_{i_1} \cap ..\hat{D_{i_j}}.. \cap D_{i_q}$ obtained by possible omitting the indices, on $A_D^{pq}$, degenerates at $E_2$, with 

\vspace{0.5cm}

\begin{center} 
$E_1^{pq}=H^q(D^{(p+1)},\mathbb{C}) \Rightarrow H^{p+q}(D,\mathbb{C})$
\end{center}

\vspace{0.5cm}

\noindent
and computes the cohomologies of $D$. Moreover, the two filtrations 

\vspace{0.5cm}

\begin{center} 
$F^pA_D^{..}=\bigoplus_{r \geq p}A_D^{r.}, \qquad W_qA_D^{..}=\bigoplus_{s \geq -q}A_D^{.s}$
\end{center}

\vspace{0.5cm}

\noindent
induce the Hodge and the weight filtrations on $H^k(D,\mathbb{C})$ for each $k$, to define a mixed Hodge structure. 

\end{lemma}

\vspace{0.5cm}

\begin{proposition} \cite{JS2}
The spectral sequence of $B^{pq}:=A^{pq}/W_q$ degenerates at $E_2$ term with

\vspace{0.5cm}

\begin{center} 
$E_1^{-r,q+r}= \bigoplus_{k\geq 0, r}H^{q-r-2k}(D^{(2k+r+1)},\mathbb{C})(-r-k) \Rightarrow H^q(X_{\infty},\mathbb{C})$
\end{center}

\vspace{0.5cm}

\noindent
and equips $H^q(X_{\infty},\mathbb{C})$ with a mixed Hodge structure.

\end{proposition}

\vspace{0.5cm}

\noindent
Thus we assume after suitable blow ups the family is given by  $f:J \to \mathbb{C}$, where afar from $0$ the fibration looks like fibers of determinant, and $J_0=f^{-1}(0)$ is a normal crossing divisor with quasi-unipotent monodromies. One can equip $H^m(J_t), \ H^m(J_0), \ H_m(J_0)$ with mixed Hodge structures. The formation of the weight and Hodge filtrations is standard in this case, and one may filter the associated double complexes naively in vertical and horizontal directions. The MHS fit in the Clemens-Schmid exact sequence (\cite{SP} page 285, \cite{ITW}) as mixed Hodge structures,

\[ ... \to H_{2n+2-m}(J_0) \stackrel{\alpha}{\rightarrow} H^m(J_0) \stackrel{i_t^*}{\rightarrow} H^m(J_t) \stackrel{N}{\rightarrow} H^m(J_t) \stackrel{\beta}{\rightarrow} H_{2n-2m}(J_0) \stackrel{\alpha}{\rightarrow} H^{m-2}(J_0) \to ... \]

\vspace{0.5cm}

\noindent
where $\alpha $ is induced by Poincare duality followed by projection, and $\beta$ is by inclusion followed by Poincare duality. The monodromy weight filtration on $H^m(J_0)$ can be described using the hypercover structure obtained by intersections of NC divisors. Then the weight filtration of $H^m(J_t)$ can be computed via the induced filtration on $\ker(N):=K_t^m$, and satisfies;

\vspace{0.3cm}

\[ Gr_k^WH^m(J_t) \cong \begin{cases}  Gr_kK_t^m \oplus Gr_{m-2}K_t^m \oplus ... \oplus Gr_{k-2[k/2]}K_t^m , \qquad k \leq m \\[0.3cm]
 Gr_{2m-k}H^m(J_t) , \qquad k >m \end{cases} \]

\vspace{0.5cm}

\noindent
The relations between the weight filtrations can be explained by the followings, cf. \cite{ITW};

\vspace{0.5cm}

\begin{itemize}
\item $i_t^*$ induces 

\[ 
Gr_kH^m(J_0) \cong Gr_kK_t^m 
\]

\vspace{0.3cm}

\item The following sequence is exact

\[ 
0 \to Gr_{m-2}K_t^{m-2} \to Gr_{m-2n-2}H_{2n+2-m}(J_0) \stackrel{\alpha}{\rightarrow} Gr_mH^m(J_0) \to Gr_mK_t^m \to 0 
\]

\vspace{0.5cm}

\end{itemize}

\vspace{0.5cm}

\noindent
Now assume $J \hookrightarrow \bar{J}$ is a projective compactification such that $J$ is Zariski dense in $\bar{J}$. Then we have the long exact sequence

\vspace{0.3cm}

\[ ...\to H^{m-1}(\bar{J}\setminus J_0) \to H^m(\bar{J},\bar{J} \setminus J_0) \to H^m(\bar{J}) \to H^m(\bar{J} \setminus J_0) \to ... \]

\vspace{0.3cm}

\noindent
where the morphisms are of MHS, and the isomorphism

\vspace{0.3cm}

\[ H^{2n+2-m}(\bar{J},\bar{J} \setminus J_0) \cong H_c^{2n+2-m}( J) \cong H^m(J)^{\vee} \]

\vspace{0.3cm}

\noindent
as MHS. In our case if we blow up $\bar{J_{\mathfrak{m}}}$ along $\Theta_{\mathfrak{m}}$
then the above method computes the mixed Hodge structure of $H^{\pi}(\bar{J} \setminus D)=H^{\pi}(\bar{J}_{\mathfrak{m}} \setminus \Theta_{\mathfrak{m}})$. 

\vspace{0.5cm}

\begin{remark} \cite{ITW}
For non-singular projective curves of genus $g$ the $H^{g-1}(\Theta,\mathbb{Z})$ contains a natural sublattice

\vspace{0.3cm}
 
\[
P=\ker (H^{g-1}(\Theta, \mathbb{Z}) \to H^{g+1}(J,\mathbb{Z}) )
\]

\vspace{0.3cm}

\noindent
where the map is the pushforward by inclusion. Then, as Hodge structures

\vspace{0.3cm}

\[ 
H^{g-1}(\Theta) =P \oplus \Theta .H^{g-3}(J)
\]

\vspace{0.3cm}

\end{remark}

\vspace{0.5cm}

In this text we apply some of the consequences of the graph tree theorem to the theory of the motives of graphs and try to relate it to the mixed Hodge structures of configuration polynomials.

\section{Motives of graphs}

\vspace{0.5cm}

Assume $\Gamma$ is a connected graph, with $h_1(\Gamma):=\text{rank} H^1(\Gamma)$, and $\sharp E(\Gamma)=m$. Then we have the homology sequence 

\[ 0 \to H_1(\Gamma,\mathbb{Z}) \to \mathbb{Z}[E(\Gamma)] \to \mathbb{Z}[V(\Gamma)] \to H_0(\Gamma,\mathbb{Z}) \to 0 \]

\vspace{0.3cm}

\noindent
$\mathbb{Z}[E(\Gamma)]$ is canonically self dual, by simply taking the dual basis to the basis given by the edges $e \in E$. For each edge $e \in E$, the functionals 
$e^{\vee}$, restrict to $H$ and thus, $(e^{\vee})^2:H \to \mathbb{C}$ defines a rank one quadric on $H$. 

\begin{definition}
The graph polynomial is defined by;

\[ \Psi_{\Gamma}:=\Psi_H=\det (\sum_e x_e.(e^{\vee})^2) \]

\end{definition}

\vspace{0.2cm}

\noindent
The definition depends only on the configuration $H \subset \mathbb{Z}[E]$ and not on the graph.

\vspace{0.2cm}

\begin{proposition} \cite{B1}
$\Psi_{\Gamma}:=\sum_T \prod_{e \notin T} x_e$, where $T$ runs over all the spanning trees $T \subset \Gamma$.
\end{proposition}

\vspace{0.2cm}

\begin{proposition} \cite{B1}
$\Psi_{\Gamma}=\prod_e x_eL(\Gamma)(x_e)$, where $L$ is the Laplacian calculated in edge weight of $e$ to be $x_e$
\end{proposition}

\vspace{0.2cm}

The motive of the graph $\Gamma$ is 

\[ M(\Gamma):=H^{m-1}(\mathbb{P}^{m-1}-X_{\Gamma}, \mathbb{Q}(m-1))  \]

\vspace{0.2cm}

\noindent
where $X_{\Gamma}:\Psi=0$ is the graph hypersurface. 

One of the classical results in graph theory is the matrix-tree theorem which asserts that the determinant of the cofactor of the combinatorial Laplacian is equal to the number of spanning trees in the graph. The usual notion of Laplacian for graphs considers edge weights. It is not obvious how to define the Laplacian with vertex weight. 

Let $G$ be a graph with vertex set $V=\{v_1,...,v_n\}$ and edge set $E$. The Laplacian of $G$ is defined by 

\[ L(u,v)=\begin{cases} d_v, \qquad u=v \\-1 \qquad \ \text{if} \  u \sim v \ \text{adjacent} \\
0 \end{cases} \]

\vspace{0.2cm}

\noindent
where $d_v$ denotes the degree of $v$. In order to generalize the definition of the Laplacian to graphs with vertex weight $\alpha_v$. Define the  matrix

\[ \mathbb{L}(u,v)=\begin{cases} \sum_{z \sim v} \alpha_v, \qquad u=v \\-\alpha_v \qquad \ \text{if} \ u \sim v \ \text{adjacent} \\
0 \end{cases} \]

\vspace{0.2cm}

\noindent
The matrix $\mathbb{L}$ is not symmetric. To make it symmetric define the following matrix which is equivalent to $\mathbb{L}$;

\[ \mathcal{L}(u,v)=\begin{cases} \sum_{z \sim v} \alpha_v, \qquad u=v \\-\sqrt{\alpha_u.\alpha_v} \qquad \text{if} \ u \sim v \ \text{adjacent} \\
0 \end{cases} , \qquad \mathbb{L}=W^{-1/2}\mathcal{L}W^{1/2} \]

\vspace{0.2cm}

\noindent
where $W$ is the $n \times n$-diagonal matrix with $(v,v)$-entry $\alpha_v$. Then $\mathcal{L}$ is symmetric. In fact if $B$ is the incidence matrix with rows indexed by vertices and columns indexed by edges defined by

\[ B(v,e)=\begin{cases} \sqrt{\alpha_v}, \qquad e={u,v}, \ u=u_i, v=v_j, i<j \\-\sqrt{\alpha_v} \qquad \text{if} \ e={u,v}, \ u=u_i, v=v_j, i>j \\
0 \end{cases} \Rightarrow \qquad \mathcal{L}=BB^* \]

\vspace{0.2cm}

\noindent
where $B^*$ is the transpose of $B$. Suppose Now the graph $G$ also has the edge weight $w_e=w_{u,v} \geq 0$, and let $T$ be the diagonal matrix with $(e,e)$-entry $w_e$. Then a result proved in \cite{CL} is 

\begin{equation} 
\mathcal{L}=BTB^* 
\end{equation}

\vspace{0.2cm}

\noindent
This equation can be regarded as a generalization for the Laplacian when vertex and edge weights. 

Let $G$ be our graph with $v$ having weight $\alpha_v$ and $T$ a tree in $G$. For $v \in T$ define the rooted directed tree $T_v$, by orienting every edge of $T$ toward the root $v$. Define the edge set of $T_v$ by

\[ E(T_v)=\{ \ (x,Y) \| \ d_T(v,x) >d_T(v,x) \ \} \]

where $d_T(v,x)$ is the distance in $T$ between $v$ and $x$. Define

\[ w(t_v)=\prod_{(x,y) \in E(T_v)}\alpha_v \qquad \text{weight of} T_v \]
\[ k_v(G)=\sum_T w(T_v) , \qquad k(G)=\sum_vk_v(G) \]

\vspace{0.2cm}

\noindent
If $\alpha_v=1$ then $k(G)$ is the number of rooted directed spanning trees in $G$. We use this definition to extend the definition of the motive defined by a graph when having weights. According to \cite{CL}, the determinant of any cofactor of $\mathcal{L}$ obtained by deleting the $u$-th row and $v$-th collumn is 

\[ \sqrt{\alpha_u. \alpha_v}(\sum_z \alpha_z)^{-1}k(G) \]

\vspace{0.2cm}

\noindent
A forest in a graph is a subgraph containing no loops. Let $S \subset V(G)$ with $|S|=s$, and $X \subset E$ with $|X|=n-s$ be subsets. If the subgraph with vertex set $V(G)$ and edge set $X$ is a spanning forest and each of its subtree contains exactly one vertex in $S$, Define the rooted directed spanning forest $X_S$, consisting of all edges of $X$ oriented toward $S$. Set

\[ w(X_S)=\prod_{(x,y) \in E(X_S)} \alpha_y , \qquad \text{weight of} \ X_S \]

\[ k_S(G)=\sum_{X_S}w(X_S) , \qquad k_S(G)=\sum_{S, \ |S|=s} k_S(G) , \ 1 \leq s \leq n \]

\vspace{0.2cm}

Then we have the following well known theorem 

\vspace{0.2cm}

\begin{theorem}\cite{CL}
The $s$-th coefficient of the characteristic polynomial of 
$\mathcal{L}$ is the sum of weights of all the rooted directed spanning forests with $s$-roots.
\end{theorem}

\vspace{0.2cm}

Along these discussions we can generalize the definition of the motive by graphs to weighted one. That is to define 

\[ \Psi_{\Gamma}:= \prod_ex_e \mathcal{L}(\alpha_v,x_e) \]

\vspace{0.2cm}

\noindent
where $\mathcal{L}(\alpha_v,x_e)$ means the Laplacian is computed with vertex weight $\alpha_v$ and edge weight $x_e$. 

\vspace{0.2cm}
 
\section{Appendix: Extensions of groups}

\vspace{0.5cm}

Let $A,B,C$ be three algebraic groups. A sequence of algebraic homomorphisms $0 \to B \to C \to A \to 0$ is called strictly exact If it is exact in the usual sense, and if all the homomorphisms are separable, In other words if the algebraic structure of $B$, and $A$ is induced by that of $C$. This would be equivalent to 

\[ 0 \to \mathfrak{t}_B \to \mathfrak{t}_C \to \mathfrak{t}_A \to 0 \]

\vspace{0.2cm}

\noindent
is exact. A strict exact sequence would be called an extension of $A$ by $B$. The set of isomorphism classes of all extensions of $A$ by $B$ is denoted $Ext(A,B)$, which is endowed with a composition law as follows. For $f:B \to B^{\prime}$ and $C \in Ext(A,B)$ then $f_*C \in Ext(A,B^{\prime})$ may be defined as the quotient of $C \times B^{\prime}$ by the subgroup formed by $(-b,f(b))$. The element $f_*C$ is the unique extension such that there exists a homomorphism $F:C \to C^{\prime}$ making a commutative diagram

\begin{equation}
\begin{CD}
0 @>>> B @>>> C @>>> A @>>> 0 \\
@|    @VVV   @VVV  @VVV  @|\\
0 @>>> B^{\prime} @>>> C^{\prime} @>>> A @>>> 0
\end{CD}.
\end{equation}

\vspace{0.2cm}

\noindent
Similarly one defines the pull back $g^*C$ for $g:A^{\prime} A$ as the subgroup of $A^{\prime} \times C$ formed by pairs having the same image in $A$. If $C,C^{\prime}$ are two elements in $Ext(A,B)$, then $C \times C^{\prime}$ can be considered as an element of $Ext(A \times A,B \times B)$, denoting by $d:A \to A \times A$ and $s:B \times B \to B$ the diagonal of $A$ and composition law of $B$, then $C+C^{\prime}=d^*s_*(C \times C^{\prime})$ in $Ext(A,B)$ and , makes it into an additive bi-functor, \cite{S} Chap VII.

\vspace{0.2cm}

\begin{remark} (\cite{S} page 166)
The category $\mathcal{C}_k$ of commutative algebraic groups over the field $k$ is an additive category. It is an Abelian category, when the characteristic of $k$ is $0$ and only in this case. However, $\mathcal{C}_k$  contains neither enough injectives nor enough projectives. 
\end{remark}

\end{document}